# Language Models are Crossword Solvers

### v2

@@ Appendix A Appendix - Details of the Sweep Clip Algorithm
A crossword puzzle consists of a grid (see Figure 1) and a set of clues $C=\{c_{1},\ldots c_{n}\}$ and answers (ground truth) $A=\{a_{1},\ldots a_{n}\}$ corresponding to position $i$ in the grid. The grid imposes a graph $G=(V,E)$, where $V=\{v_{1},\ldots v_{n}\}$ is the set of vertices corresponding to every clue/answer in the crossword, and
@@
Given an LLM, and a set of clues $C^{\prime}\subseteq C,C^{\prime}=\{c_{j_{1}},c_{j_{2}},\ldots\}$ corresponding to vertices $\{v_{j_{1}},v_{j_{2}},\ldots\}$ we can generate candidate answers $\hat{A}=\{\hat{a}_{j_{1}},\hat{a}_{j_{2}},\ldots\}$, where $\hat{a}=LLM(c)$. We abbreviate this as $\hat{A}=LLM(C^{\prime})$.
@@
There are two sub-graphs $G_{p}(\hat{A})$ and $G_{n}(\hat{A})$ of $G$ that correspond to a set of candidate answers $\hat{A}$. Let $v_{i},v_{j}$ correspond to $\hat{a}_{i},\hat{a}_{j}\in\hat{A}$. An edge $(v_{i},v_{j})\in E$ is in $G_{p}(\hat{A})$ if and only if $\hat{a}_{i},\hat{a}_{j}$ don’t conflict, else it is in $G_{n}(\hat{A})$. We call the largest connected component of a graph $H$, $LCC(H)$, and the neighborhood of a set of vertices $V$ in the graph $G$, $ngbd(G,V)$. The algorithm is detailed in Algorithm 2.
@@
12:     Calculate $N=ngbd(G,\hat{A})$.
13:     $\hat{A}=\hat{A}+LLM(N)$
14:     if solved or budget_exceeded then
15:         break
16:     end if
17:end for
@@ Appendix B Appendix - Datasets
Dataset (Clues) Train Validation Test
@@
Table 7: Dataset details
@@ B.1 Cryptonite and word-init-disjoint
Cryptonite is a cryptic crossword clue dataset introduced by Efrat et al. (2021), and is gathered from The Times and The Telegraph - two popular UK dailies. In their official split, the training and test sets do not contain any common answers, so that a model trained on the train split cannot rely on memorization to answer the cryptic crossword clue. Two thousand random samples samples were chosen from the train split to report results in Section 3. The Cryptonite dataset [^13] https://github.com/aviaefrat/cryptonite is not distributed under an open source licence.
word-init-disjoint is also a cryptic crossword clue and was introduced by Rozner et al. (2021). This is gathered from The Guardian, and their official split introduces the additional constraint that if two answers share a common root (first two characters) they are placed in the same split. The results reported in 3 are based on two thousand randomly sampled clues from their official test split. We distribute all materials pertaining to this dataset [^14] https://github.com/jsrozner/decrypt with this paper.
@@ B.2 New York Times Dataset
The test and support sets are distributed alongside this paper in the Supplementary Materials, however, the 100 complete grids are not, as they are intellectual property of the New York Times. They can be accessed after purchasing a subscription to their service. To aid reproducibility we include the list of dates to uniquely identify the crossword grids used to report results.
@@ B.3 Post-cutoff datasets
We curated two post-cutoff datasets to test for potential dataset contamination. One of them was sourced from Lovatts puzzles  [^17] https://lovattspuzzles.com/ and The Daily Mail. In total 244 clues and answers were manually curated from puzzles published strictly after May 20, 2024. To maintain parity with word-init-disjoint dataset by Rozner et al. (2021) we also obtained 200 clue answer pairs from The Guardian puzzles published after May 20, 2024.
We cross-checked both these datasets for duplicate clues against all 665,497 clues from the Cryptonite and word-init-disjoint datasets. We found two duplicates for the Lovatts set which were removed. The dataset obtained from The Guardian contained no duplicates. This does not guarantee that these cryptic crossword clues have not appeared in any form previously, but we estimate that probability to be low, since there were very few duplicates when matched against the extremely large cryptic crossword datasets. Both these datasets are distributed alongside the paper in the Supplementary Materials.
@@ B.4 Sub-token counting
For the task in Section 3.4 we first obtained a list of 300,000+ words and their unigram frequencies provided by Segaran and Hammerbacher (2009) and sourced from the Google Trillion Words corpus. For the Rare, Medium and Common words we randomly selected 1000 words each from ranks 95,000-100,000, 47,500-52500 and 0-5000 respectively. The Vocab set was constructed by taking an intersection of the vocabulary tokens of every open source language model in consideration, i.e. Phi 3, LLaMA 3 70B, LLaMA 3 8B, LLaMA 2, Mistral 7B, and Mixtral. Further, we took the intersection of this set with the set of top 100,000 words to make sure the chosen tokens are actual words and not special tokens. We sampled 1000 words from the resulting set to create the vocab set. For the gibberish set, we replaced each character from each word in every set with a character randomly sampled from the set {a-z} and randomly sampled 1000 words from it. This ensures that the distribution of lengths of words seen in the gibberish set is close to the distribution of lengths in the other sets. These 5 sets are distributed alongside the paper in the Supplementary Materials.
@@ C.1 General Overview
We used temperature=0.5 for all models and experiments. For few-shot prompt responses max-tokens were set to 10, and for Chain-of-thought responses max-tokens was set to 1000. top_p, top_k or any other parameters were not changed. The Jinja templates used to format the text prompts are provided alongside the paper, and are borrowed from the model-cards of the respective models. Claude and GPT models were accessed through their respective APIs and all other models were run locally on one server consisting of 2x80GB Nvidia A100 GPUs. All models were used in bf16 format whenever supported. Format of prompts are given in Appendix D.
@@ Appendix D Appendix - Prompts
In this section we will detail the prompts used for all experiments presented in the paper. Every prompt used in generating outputs presented in this paper is provided alongside the paper in Supplementary Materials.
@@ D.1 Few Shot Prompts
For the experiments presented in Section 3.2 we used prompts in the following format -
@@
For the experiments presented in Section 3.3 we used prompts in the following format (with the same system prompt) -
@@ D.2 Chain-of-thought Prompts
For the results presented in Table 3 and for human-evaluation we used Chain-of-thought prompts. These were based on 10 examples we manually created and are diverse in terms of skills required to solve them. An example Chain-of-thought reasoning prompt is as follows -
@@ Appendix E Appendix - Human Evaluation and Reasoning Examples
Human evaluation was performed by 3 of the authors of this paper. They were asked if a particular response contained logical or factual errors and provided detailed instructions and examples of the kinds of errors to be wary of. All responses were evaluated by at least 2 authors and if there were conflicting answers, discussions were held to reach a consensus. All evaluated responses alongside human-annotated labels are provided alongside this paper.
In this section we provide a few Chain-of-thought elicited reasoning examples given by GPT-4-Turbo on the Lovatts dataset. These are sampled from the set used to perform human evaluation.
@@ Appendix F Appendix - Results
Table 9: Data for results plotted in Figure 3
@@
Table 10: Data for results plotted in Figure 2

### v3

@@ Appendix A Appendix - Details of the Sweep Clip Algorithm
A crossword puzzle consists of a grid (see Figure 1) and a set of clues $C=\{c_{1},\ldots,c_{n}\}$ and answers (ground truth) $A=\{a_{1},\ldots,a_{n}\}$. The grid imposes a graph $G=(V,E)$, where $V=\{v_{1},\ldots,v_{n}\}$ is the set of vertices corresponding to every clue/answer in the crossword, and
@@
Given an LLM, and a set of clues $C^{\prime}\subseteq C,C^{\prime}=\{c_{j_{1}},c_{j_{2}},\ldots\}$ corresponding to vertices $\{v_{j_{1}},v_{j_{2}},\ldots\}$, we can generate candidate answers $\hat{A}=\{\hat{a}_{j_{1}},\hat{a}_{j_{2}},\ldots\}$, where $\hat{a}=LLM(c)$. We abbreviate this as $\hat{A}=LLM(C^{\prime})$. For a subset $V^{\prime}$ of $V$, let the set of clues associated with $V$ be denoted by $C(V^{\prime})$.
In subsequent iterations of the algorithms, when some candidate answers have been generated, we generate further candidate answers for their neighboring vertices. In this case, the partially unmasked characters are provided to the LLM as in Section 3.3. Note, our algorithm implicitly performs self-consistency checks to improve candidate answers, e.g., an answer generated at the first sweep may be discarded only to be accepted in later iterations, when it is coherent with a larger number of other answers.
@@
There are two sub-graphs $G_{p}(\hat{A})$ and $G_{n}(\hat{A})$ of $G$ that correspond to a set of candidate answers $\hat{A}$. Let $v_{i},v_{j}$ correspond to $\hat{a}_{i},\hat{a}_{j}\in\hat{A}$. An edge $(v_{i},v_{j})\in E$ is in $G_{p}(\hat{A})$ if and only if $\hat{a}_{i},\hat{a}_{j}$ don’t conflict, else it is in $G_{n}(\hat{A})$. We call the largest connected component of a graph $H$, $LCC(H)$, and for a subset $S$ of vertices $V$ of a graph $G$, $ngbd(G,S)$ denotes the vertices in $V$ that are adjacent to $S$ but not in $S$. The algorithm is detailed in Algorithm 1.
@@
12:     Calculate $N\leftarrow ngbd(G,\hat{A})$.
13:     Add character information to $C(N)$.
14:     $\hat{A}\leftarrow\hat{A}\cup LLM(C(N))$
15:     if solved or budget_exceeded then
16:         break
17:     end if
18:end for
The results in Section 5.1 are produced with max_iter of 30 and a budget of 0.5 USD per crossword for GPT-4-Turbo, and max_iter of 35 and a budget of 600 LLM calls for LLaMA 3 70B.
@@ Appendix B Appendix - Datasets
Dataset Train Validation Test
@@
Table 8: Dataset details.
@@ B.1 Cryptonite and word-init-disjoint
Cryptonite [^13] https://github.com/aviaefrat/cryptonite is a cryptic crossword clue dataset introduced by Efrat et al. (2021), and is gathered from The Times and The Telegraph - two popular UK dailies. In their official split, the training and test sets do not contain any common answers, so that a model trained on the train split cannot rely on memorization to answer the cryptic crossword clue. Two thousand random samples were chosen from the train split to report results in Section 3.
word-init-disjoint (Init) is also a cryptic crossword clue dataset and was introduced by Rozner et al. (2021). This is gathered from The Guardian, and their official split introduces the additional constraint that if two answers share a common root (first two characters) they are placed in the same split. The results reported in 3 are based on two thousand randomly sampled clues from their official test split. We distribute all materials pertaining to this dataset [^14] https://github.com/jsrozner/decrypt with this paper.
@@ B.2 New York Times Dataset
The test and support sets are distributed alongside this paper in the github repository, however, the 100 complete grids are not, as they are intellectual property of the New York Times. They can be accessed after purchasing a subscription to their service. To aid reproducibility we include the list of dates to uniquely identify the crossword grids used to report results.
@@ B.3 Post-cutoff datasets
We curated two post-cutoff datasets to test for potential dataset contamination. One of them was sourced from Lovatts puzzles  [^17] https://lovattspuzzles.com/ and The Daily Mail. In total, 244 clues and answers were manually curated from puzzles published strictly after May 20, 2024. To maintain parity with the Init dataset by Rozner et al. (2021) we also obtained 200 clue-answer pairs from The Guardian published after May 20, 2024.
We cross-checked both these datasets for duplicate clues against all 665,497 clues from the Cryptonite and Init datasets. We found two duplicates for the Lovatts set which were removed. The dataset obtained from The Guardian contained no duplicates. This does not guarantee that these cryptic crossword clues have not appeared in any form previously, but we estimate that probability to be low, since there were very few duplicates when matched against the extremely large cryptic crossword datasets. Both these datasets are distributed alongside the paper in the github repository.
@@ B.4 Sub-token counting
For the task in Section 3.4 we first obtained a list of 300,000+ words and their unigram frequencies provided by Segaran and Hammerbacher (2009) and sourced from the Google Trillion Words corpus. For the Rare, Medium and Common words we randomly selected 1000 words each from ranks 95,000 - 100,000, 47,500 - 52,500 and 0 - 5,000, respectively. The Vocab set was constructed by taking an intersection of the vocabulary tokens of every open source language model in consideration, i.e., Phi 3, LLaMA 3 70B, LLaMA 3 8B, LLaMA 2, Mistral 7B, and Mixtral. Further, we took the intersection of this set with the set of top 100,000 words to make sure the chosen tokens are actual words and not special tokens. We sampled 1000 words from the resulting set to create the vocab set. For the gibberish set, we replaced each character from each word in the vocab set with a character randomly sampled from the set {a-z}. This ensures that the distribution of lengths of words seen in the gibberish set is the same as the distribution of lengths in the vocab sets. These 5 sets are distributed alongside the paper in the github repository.
@@ C.1 General Overview
We used temperature=0.5 for all models and experiments. For few-shot prompt responses, max-tokens were set to 10, and for chain-of-thought responses, max-tokens was set to 1000. All other paramters including top_p, top_k, etc., were set to their defaults. The Jinja templates used to format the text prompts are provided alongside the paper, and are borrowed from the model-cards of the respective models. Claude and GPT models were accessed through their respective APIs and all other models were run locally on one server consisting of 2x80GB Nvidia A100 GPUs. All models were used in bf16 format whenever supported. Format of prompts are given in Appendix D.
@@ Appendix D Appendix - Prompts
In this section we will detail the prompts used for all experiments presented in the paper. Every prompt used in generating outputs presented in this paper is provided alongside the paper in the github repository.
@@ D.1 Few Shot Prompts
For the experiments presented in Section 3.2, we used prompts in the following format:
@@
For the experiments presented in Section 3.3, we used prompts in the following format (with the same system prompt):
@@ D.2 Chain-of-Thought Prompts
For the results presented in Table 4 and for human-evaluation we used chain-of-thought prompts. These were based on 10 examples we manually created and are diverse in terms of skills required to solve them. An example chain-of-thought reasoning prompt is as follows:
@@ Appendix E Appendix - Human Evaluation and Reasoning Examples
Consider the following example cryptic crossword clue - ‘‘Culminating point of story about Judy’s husband by the railway track (5,4)’’.  [^18] Example taken from https://lovattspuzzles.com. Solving this clue requires connecting ‘Judy’ to the popular puppet show Punch and Judy, and inferring that ‘Judy’s husband’ refers to PUNCH. Additionally, we must observe that ‘railway track’ is synonymous to LINE, and combining these gives PUNCHLINE which also means ‘Culminating point of story’. In this section we provide a few Chain-of-thought elicited reasoning examples given by GPT-4-Turbo on the Lovatts dataset. These are sampled from the set used to perform human evaluation.
Further, these samples are manually tagged with the principal skill required to solve them. The skill-based categories are:
  • ANG - the answer is an anagram of some words of the clue (e.g., CUBIT IS mixed up cookie $\rightarrow$ BISCUIT).
  • HOM - the answer is a homophone of some words of the clue (e.g., Heard PRINTS are for royalty $\rightarrow$ PRINCE).
  • CNT - the answer is disguised in a contiguous section of the clue (e.g., The Press leaves presENTERs to go in $\rightarrow$ ENTER).
  • SCJ - the answer is found by combining several words which are synonyms of various parts of the clue (e.g., Reasonable FOOD ALLOWANCE1 for CAPONE2 $\rightarrow$ RATION1AL2).
  • OTH - this class lumps together variety of other kinds of skills like Spoonerisms, acronyms, world knowledge and various other kinds of character manipulations (e.g. PIpER loses heart on jetty $\rightarrow$ PIER).
Human evaluation was performed by 3 of the authors of this paper. They were asked if a particular response contained logical or factual errors and provided detailed instructions and examples of the kinds of errors to be wary of. All responses were evaluated by all authors and in case of conflicting answers (4 out of 100, Fleiss’ $\kappa$ = 0.94), discussions were held to reach a consensus. All evaluated responses alongside human-annotated labels are provided alongside this paper in the github repository.
In checking whether a justification for an answer given by the LLM is logically and factually sound, we assess grammatical soundness and phraseological meaningfulness of the sentences in the answer, existence of counterfactual statements (e.g., “BULKY has 4 characters”, “the initial letters of ARE RATS TIRED NOW are ARTS” etc.), presenting a statement as an inference from previous statements when it does not follow from those, etc. If an answer by the LLM is found unsatisfactory in any of these aforementioned areas, it is labeled FALSE. Following are a few samples of the same.
@@ Appendix F Appendix - Results
Table 10: Data for results plotted in Figure 3.
@@
Table 11: Data for results plotted in Figure 2.